Given the following cost function: C(q) = - 8+30q+5, find AC


1
Expert's answer
2021-09-09T18:11:07-0400

 Cost function: "C(q) = -8 + 30 q + 5"

Average Cost: "AC = \frac{C(q)}{q}"

"AC = \frac{-8}{q} +\frac{30q}{q} + \frac{5}{q}"

"AC = \frac{-8}{q} + 30 + \frac{5}{q}"